OpenStudy (anonymous):

cosx+2secx=-3 solve each equation for all real values of x

OpenStudy (mertsj):

change secx to 1/cosx

OpenStudy (mertsj):

\[\cos x+\frac{2}{\cos x}+3=0\]

OpenStudy (mertsj):

\[\cos ^2x+3\cos x+2=0\]

OpenStudy (mertsj):

Factor and solve.

OpenStudy (anonymous):

could you elaborate a little on how you got from the 2/cosx to the next step?

OpenStudy (anonymous):

nevermind! thanks